Marula Enterprises has developed a new local brew called Stiff Neck. The market demand for this product is given as follows: Q=240-4p. If Stiff Neck is priced at R40, what is the point elasticity of demand?
1
Expert's answer
2020-03-19T11:57:41-0400

P=40 Q=80

"\Delta"Q/"\Delta"P=-4

e = -4*P/Q = -4*40/80 = -2